Hi! My teacher recently gave us a worksheet that requires us to use systems of equations, I was hoping you'd be able to help me. Thank you! :D
If you add five to the sum of Patrick's age and his mother's age, the total is 48. Patrick's mother is five years less than three times the age of Patrick. How old are Patrick and his mother?
Thank you! :)

Answer by jim_thompson5910(35256)

Let
P = Patrick's age
M = Mother's age


"If you add five to the sum of Patrick's age and his mother's age, the total is 48", so we know that 5+P+M = 48

5+P+M = 48

P+M = 48-5

P+M = 43

M = 43-P

-------------------------------------------
"Patrick's mother is five years less than three times the age of Patrick", so we also know that

M = 3P - 5

43 - P = 3P - 5 ... plug in M = 43 - P

I'll let you finish up. You would solve this last equation for P. Once you have P, you can find M.